Equivalent Units of Production The following information concerns production in the Baking Department for December. All...

  1. Equivalent Units of Production

    The following information concerns production in the Baking Department for December. All direct materials are placed in process at the beginning of production.

    ACCOUNT Work in Process—Baking Department ACCOUNT NO.
    Date Item Debit Credit Balance
    Debit Credit
    Dec. 1 Bal., 6,000 units, 3/5 completed 8,640
    31 Direct materials, 108,000 units 194,400 203,040
    31 Direct labor 56,060 259,100
    31 Factory overhead 31,540 290,640
    31 Goods finished, 109,500 units 279,660 10,980
    31 Bal., ? units, 4/5 completed 10,980

    a. Determine the number of units in work in process inventory at December 31.
    units

    b. Determine the equivalent units of production for direct materials and conversion costs in December. If an amount is zero, enter in "0".

    Baking Department
    Equivalent Units of Production for Direct Materials and Conversion Costs
    For December
    Whole Units Equivalent Units
    Direct Materials
    Equivalent
    Units Conversion
    Inventory in process, December 1
    Started and completed in December
    Transferred to finished goods in December
    Inventory in process, December 31
    Total
Answer #1

a. Determine the number of units in work in process inventory at December 31 units

Unit in work in process, January 1 6000
Add: Units started during the year 108000
Less: Units completed and transferred (109500)
Units in work in process, December 31 4500

b. Determine the equivalent units of production for direct materials and conversion costs in December. If an amount is zero, enter in "0".

Baking Department
Equivalent Units of Production for Direct Materials and Conversion Costs
For December
Whole Units Equivalent Units
Direct Materials
Equivalent
Units Conversion
Inventory in process, December 1 6000 0 (6000*2/5)= 2400
Started and completed in December 103500 103500 103500
Transferred to finished goods in December 109500 103500 105900
Inventory in process, December 31 4500 4500 (4500*4/5)= 3600
Total 114000 108000 109500

Units started and completed= Units completed-Beginning units in work in process

= 109500-6000= 103500 units